OpenStudy (anonymous):

Help? It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? How do I solve this?

OpenStudy (anonymous):

nvm sorry i misread

OpenStudy (anonymous):

let the distance be x average speed = total distance/total time = 2x/12 = 20 mph therefore x = 120 miles

OpenStudy (anonymous):

i think i read wrong total time is not 12 hrs.

OpenStudy (anonymous):

So if he can go 20 mph out, that's 6 hours out... and 30 mph back, that's 4 hours back?

OpenStudy (anonymous):

well, no, because that's only 7 hrs..

OpenStudy (anonymous):

It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? : Let d = dist traveled one way Write a time equation; time = dist/speed d/20+d/30 = 12 multiply eq by 60, results 3d + 2d = 60(12) 5d = 720 d = 720/5 = 144 mi one way : Find the time at each speed 144/20= 7.2 hrs at 20 mph 144/30 =4.8 hrs at 30 mph Does this work?

OpenStudy (anonymous):

That helps a lot!

OpenStudy (anonymous):

distance coming back = 30mph x 12 hrs = 360 miles

OpenStudy (anonymous):

I am not going to plagerize....here is the site i found it from http://www.algebra.com/algebra/homework/word/travel/Travel_Word_Problems.faq.question.315526.html

OpenStudy (anonymous):

I needed a refresh in this anyways...thnx

OpenStudy (anonymous):

Thanks for the integrity @Mpost1994

OpenStudy (anonymous):

No problem